The probability that a football player will play Eden garden is 0.6 and on Ambedkar Stadium is 0.4. The probability that he will get knee injury when playing in Eden is 0.07 and that in Ambedkar stadium is 0.04.What is the probability that he would get a knee injury if he played in Eden.

Grade:

1 Answers

Chetan Mandayam Nayakar
13 years ago

the problem seems to be wrongly worded. i take it as "if the player got a knee injury,what is the probability that the injury happened at Eden?"

from Baes' rule the answer is (0.6*0.07)/((0.6*0.07)+(0.4*0.04))=0.042/0.058 =42/58=21/29

'Win exciting gifts by answering the questions on Discussion Forum.'